solve by the substitution method
3x+5y=21
x=51-9y

Answer by checkley77(12844)

3x+5y=21
x=51-9y
3(51-9y)+5y=21
153-27y+5y=21
-22y=21-153
-22y=-142
y=-132/-22
y=6 ans.
x=51-9*6
x=51-54
x=-3 ans.
3*-3+5*6=21
-9+30=21
21=21